one number is 3 more than the other. Their sum is 7. Find the numbers.
Answer by Cromlix(4381)

Other number = x
One number = x + 3
x + (x + 3) = 7
2x + 3 = 7
2x = 7 - 3
2x = 4
x = 2
Other number = 2
One number = 5
Hope this helps.
:-)